OpenStudy (anonymous):

Write a polynomial function f of least degree that has rational coefficients , a leading coefficient of 1 and th given zeros . 1.) -7 , -4 2.) 1,2,5

OpenStudy (campbell_st):

is a is a zero then (x - a) is a factor... you need this idea to find the functions so is simply a case for 1. f(x) = (x + 7)(x + 4) 2.f(x) = (x -1)(x -2)(x -5) just distribute to get the polynomials...